# Anomalous Change Point Detection Using Probabilistic Predictive Coding

### v1

@@ 3.4 Training procedure
$\mathcal{L}_{\text{MLE}}$ $=-\frac{1}{N_{f}}\sum^{N_{f}}_{i=1}\log f\left(z_{N_{p}+i}|\hat{z}_{N_{p}+i}\right)$
@@
where $\theta$ is the set of model parameters for models $E$, $G$, $F_{0}$, $F_{1}$, $\cdots$, $F_{N_{f}}$, and $D$, and $\lambda$ is the weight assigned to the mean squared error. After training, the decoder $D$ is no longer needed for the anomalous change point detection and can be discarded.
@@ 3.5 Probability of conformance
Now, if we consider the relationship with the probability of conformity in the data space, $p_{x,N_{p}+i}$, we must note that the relationship between $f\left(z_{N_{p}+i}|\hat{z}_{N_{p}+i},\sigma_{N_{p}+i}\right)$ and $p_{z,N_{p}+i}$ is strictly monotonically increasing. The same can trivially be said about the relationship between $f\left(x_{N_{p}+i}|x_{1},x_{2},\cdots,x_{N_{p}}\right)$ and $p_{x,N_{p}+i}$. Due to the proportionality between $f\left(z_{N_{p}+i}|\hat{z}_{N_{p}+i},\sigma_{N_{p}+i}\right)$ and $f\left(x_{N_{p}+i}|x_{1},x_{2},\cdots,x_{N_{p}}\right)$ (see section 3.3), we can therefore conclude that the relationship between $p_{z,N_{p}+i}$ and $p_{x,N_{p}+i}$ is also strictly monotonically increasing. This indicates that the ordering of values of $p_{x,N_{p}+i}$ is preserved in its proxy $p_{z,N_{p}+i}$, making this estimated probability of conformity in the latent space a useful metric for anomaly detection.
@@ 4.1 Proportionality test
The pipeline consists of an encoder with a single fully connected layer to a latent space encoding with a size of $N_{e}=4$ and a decoder with a fully connected layer back to a single value. A gated recurrent unit (GRU) model with $N_{g}=8$ units is used for predictions together with a forecasting model consisting of 8 fully connected layers with $[16,16,32,32,64,64]$ units, followed by two separate fully connected layers for the estimation of $\hat{z}_{2}$ (linear activation) and $\sigma_{2}$ (exponential activation). Training is performed using the RMSprop optimizer ($lr=1\cdot 10^{-4}$, $\rho=0.9$) with a batch size of 64. The models are optimized for 1000 warm-up iterations based on the loss in Equation 4, after which training with the normal loss function (Equation 3) continues until convergence. The loss weight for the reconstruction loss is set to $\lambda=100$. Training is repeated 100 times to determine sensitivity to initial conditions, resulting in 100 models.
@@ 4.2 Sine wave frequency deviation
$f_{c}(t)$ =$=\begin{cases}f_{c,before}&\text{if }t\text{ before change point}\\ f_{c,after}&\text{if }t\text{ after change point}\end{cases}.$ (10)
A signal containing the instantaneous frequencies at every time step $f(t)$ is generated using a constrained random walk, such that $f_{c}(t)-\frac{f_{b}}{2}\leq f(t)\leq f_{c}(t)+\frac{f_{b}}{2}$, where $f_{b}=0.25$ Hz is the frequency band. Furthermore, bounded random walks were used to create a baseline wander $-1\leq b(t)\leq 1$ and wave amplitude $0.5\leq a(t)\leq 2$. Lastly, additive noise was $n(t)\sim\mathcal{N}\left(0,\sigma_{n}^{2}\right)$, where $\sigma_{n}\in[0,0.2]$. The corresponding sine wave signal was then determined by
@@
Using this data model, we synthesized a training and validation dataset of 1,000,000 and 100,000 signals respectively at a sampling frequency of $f_{s}=128$ Hz. The signals consist of 8 segments of 256 samples each, totaling 2048 samples (16 seconds). The first $5$ segments $x_{1}$ to $x_{5}$ are used to predict the latent space representations of the last $3$ segments $x_{6}$ to $x_{8}$. The center frequencies were set such that $f_{c,before}=f_{c,after}$, implying the absence of change points. At the top of Figure 3, an example of such a non-anomalous signal is shown.
We used an encoder model $E$ with two blocks (convolutional, ReLU, BatchNorm, Maxpooling) with an increasing amount of filters (32 and 64), followed by a fully connected layer to a latent space with size $N_{e}=16$. The accompanying decoder $D$ follows a similar structure with a fully connected layer, followed by two blocks (convolutional, ReLU, upsampling) and one final convolutional layer. As the recurrent neural network $G$, we used a GRU with 32 units. The three forecasting models $F_{1}$, $F_{2}$, and $F_{3}$ consist of three fully connected layers (64, 128, and 256 units), followed by two separate fully connected layers for the estimation of $\hat{z}_{6}$ to $\hat{z}_{8}$ (linear activation) and $\sigma_{6}$ to $\sigma_{8}$ (exponential activation). All models were optimized simultaneously using the RMSprop optimizer with a learning rate of $1\cdot 10^{-4}$ and a batch size of 32. Training consisted of 1000 warm-up iterations, followed by normal training until convergence. The reconstruction loss weight was set to $\lambda=10^{4}$.
After training, the pipeline was applied to the test dataset, which was generated in a similar manner as the training and validation set, except that a change point was introduced in segment $x_{6}$. We generated the test dataset by synthesizing 10 signals for every combination $\left(f_{c,before},f_{c,after}\right)$, with a resolution of $0.05$ Hz. This results in $\left(\frac{10-0.5}{0.05}\right)^{2}\cdot 10=361,000$ test signals. In the center and lower plots of Figure 3, some example signals are shown that have such a sudden change point.
@@
Figure 3: Examples of generated signals. The top signal has no frequency deviation. The center signal has a frequency change of $1$ Hz at the change point around $t=11$ s. The bottom signal has a frequency change of $2$ Hz at the change point around $t=11$ s.
Based on this test set, we calculated the log-likelihood and probability of conformance for every pair of center frequencies $\left(f_{c,before},f_{c,after}\right)$. These metrics were averaged and are shown in Figure 4, together with the frequency band with which the training and validation signals were generated. As expected, the probability of conformance is high if $f_{c,before}\approx f_{c,after}$, and low for other combinations of center frequencies. This indicates that the model is able to detect the generated change points.
Figure 4: Average log-likelihood (left) and probability of conformity (right) per combination of center frequencies. The frequency band margins indicate how much frequencies may fluctuate as part of the generated instantaneous frequencies $f(t)$.
To test the discriminative performance of the pipeline, i.e., the ability of the pipeline to detect a sudden significant change in center frequency, we labeled all pairs $\left(f_{c,before},f_{c,after}\right)$ where $f_{c,before}=f_{c,after}$ as being true, and other pairs as false. We constructed the receiver operating characteristic (ROC) curve (see Figure 5) and determined the area under the ROC curve (AUC) to be $\mathrm{AUC}=0.99$, indicating that the pipeline has a strong discriminative ability.
@@
Figure 5: Receiver operating characteristic curve for the application of the PPC pipeline on the sine wave frequency deviation example.
@@ 4.3 Counting with MNIST digits
In this experiment, we used the Modified National Institute of Standards and Technology (MNIST) dataset [33], consisting of 60,000 training images and 10,000 test images of handwritten digits. The objective within this experiment for the pipeline is to determine whether the corresponding digits of a set of two MNIST images are sequential, e.g., if $x_{1}$ formed an image of the digit 3, $x_{2}$ should represent the digit 4. In the special case of digit 9, we considered 0 to be its successor. The experiment was designed to determine if the pipeline is able to capture (deviations from) patterns in relatively complex sequential image data.
@@
First, 10,000 validation images were selected from the original set of training images, reflecting the class composition of the training set. All image pixels were normalized to the range $[0,1]$. Data augmentation is applied online, consisting of random rotation (between $-18^{\circ}$ and $18^{\circ}$), random rectangular zooming (between $-5\%$ and $5\%$), and random translation (between $-5\%$ and $5\%$).
@@
We used an encoder model $E$ with two blocks (convolutional, ReLU, BatchNorm, Maxpooling) with an increasing amount of filters (32 and 64), followed by a fully connected layer to a latent space with size $N_{e}=16$. The accompanying decoder $D$ follows a similar structure with a fully connected layer, followed by two blocks (convolutional, ReLU, upsampling) and one final convolutional layer. As the recurrent neural network $G$, we used a GRU with 32 units. The single forecasting model $F_{1}$ consists of three fully connected layers (64, 128, and 256 units), followed by two separate fully connected layers for the estimation of $\hat{z}_{2}$ (linear activation) and $\sigma_{2}$ (exponential activation). All models were optimized simultaneously using the RMSprop optimizer with a learning rate of $1\cdot 10^{-4}$ and a batch size of 32. Training consisted of 1000 warm-up iterations, followed by normal training until convergence. The reconstruction loss weight was set to $\lambda=10^{4}$.
@@
After training, the pipeline was used to compare every pair of two images in the test set, yielding a log-likelihood and probability of conformance for every pair. These metrics were averaged over every pair of digits and are shown in Figure 6. As expected, the log-likelihoods and probabilities of conformity on the subdiagonal and top right are significantly higher than for other combinations of digits. This indicates that the model indeed is able to differentiate between successive digits and non-successive digits. Some combinations, like the combination for $(3,9)$, show higher values due to the similarity between digits, e.g., a handwritten 4 is similar in shape to the digit 9.
@@
Figure 6: Average log-likelihood (left) and probability of conformity (right) per combination of digits.
To test the discriminative performance of the pipeline, i.e., the ability of the pipeline to determine whether a pair of digits is successive, we constructed the receiver operating characteristic (ROC) curve (see Figure 7) and determined the area under the ROC curve to be $\mathrm{AUC}=0.95$. Such a high AUC indicates that the pipeline has a strong discriminative ability.
@@
Figure 7: Receiver operating characteristic curve for the application of the PPC pipeline on the successive MNIST digits example.
@@ 4.4 Artifact detection for MRSI
After training, the pipeline was used to predict all segments of every spectrum of an example in-vivo 2D MRSI slice. We further calculate the likelihoods and probabilities of conformance for every segment. Figure 9 depicts the probabilities of conformance for six example spectra located at the indicated locations as well as a probability map with pixels representing averaged spectra. Moving from the top left to top right spectrum: spectra within the homogeneous regions of the brain appear normal and show the highest probabilities of conformance, measurements near the ventricles are distorted completely, and susceptibility artifacts coming from the nasal cavity are present and detected by the method. From bottom left to right: non-anomalous spectra are consistently labeled with high probabilities, spectra close to the skull show residual lipid signal artifacts and lower probabilities of conformity, and the spectra outside the brain mask are nulled leading to no anomalies.
Figure 9: In-vivo 2D MRSI slice with indicated locations of selected segmented example spectra. The probability of conformity is shown for each segment of the spectra as well as averaged for each pixel (representing a single voxel).
@@
Figure 10: 2D brain maps for the estimated likelihoods, the metabolite concentration map for total N-acetylaspartate (NAA + NAAG) estimated by LCModel [52], as well as the corresponding CRLB(left to right).
@@ 5 Discussion
In contrast to some other methods mentioned in section 2, anomaly detection with the proposed PPC method has a linear time complexity due to the deep learning approach. This allows for scalability to large datasets. While training the PPC pipeline can be computationally expensive compared to non-neural network approaches, the models used in our experiments are remarkably lightweight compared to other deep neural network architectures. The models used in sections 4.2, 4.3, and 4.4 have $5.8\cdot 10^{5}$, $2.3\cdot 10^{5}$ and $9.5\cdot 10^{5}$ trainable parameters, respectively, while modern deep neural networks have significantly more parameters. For example, Inceptionv3 [59], ResNet-50 [20], and VGG-16 [57] have about $2.4\cdot 10^{7}$, $2.5\cdot 10^{7}$, and $1.4\cdot 10^{8}$ trainable parameters, respectively. Furthermore, where other methods may fail to provide meaningful or intuitive anomaly scores, PPC is able to provide an anomaly score that is not only normalized to a closed unit interval $[0,1]$, but also indicates a probability of data conforming to what was expected.
@@
The artifact detection for MRSI experiment (section 4.4) shows the real-world application of the method to MRSI data. The probability of conformance delivers a valuable measure for the quality of MRSI spectra and enables the filtering of artifact-ridden spectra by setting a specific threshold. This is particularly useful with high‐resolution whole‐brain acquisition techniques where hundreds of thousands of spectra are acquired while only good-quality spectra can be quantified reliably [38, 39]. Note the CRLBis only obtained after computationally intensive least-squares fitting of the spectra. In addition, having a segment-wise quality measure for each spectrum can allow metabolite-targeted filtering to improve studies of specific brain pathologies [5, 17, 50].
@@
Lastly, we recognize the lack of benchmark evaluations in this work. The lack of available real-world datasets forms an obstacle to the development and testing of CPD and AD methods [46]. Publicly accessible datasets in the field of CPD and AD are either low-complexity univariate time series, outdated, or unannotated. It is therefore no surprise that the review of [40] shows the use of a plethora of different datasets in different works, for which a significant part of them is synthetic or simulated.

### v2

@@ 3.4 Training procedure
$\mathcal{L}_{\text{MLE}}$ $=-\frac{1}{N_{f}}\sum^{N_{f}}_{i=1}\log f\left(z_{N_{p}+i}|\hat{z}_{N_{p}+i},\sigma_{N_{p}+i}\right)$
@@
where $\theta$ is the set of model parameters for models $E$, $G$, $F_{1}$, $F_{2}$, $\cdots$, $F_{N_{f}}$, and $D$, and $\lambda$ is the weight assigned to the mean squared error. After training, the decoder $D$ is no longer needed for the anomalous change point detection and can be discarded.
@@ 3.5 Probability of conformance
Now, if we consider the relationship with the probability of conformance in the data space, $p_{x,N_{p}+i}$, we must note that the relationship between $f\left(z_{N_{p}+i}|\hat{z}_{N_{p}+i},\sigma_{N_{p}+i}\right)$ and $p_{z,N_{p}+i}$ is strictly monotonically increasing. The same can trivially be said about the relationship between $f\left(x_{N_{p}+i}|x_{1},x_{2},\cdots,x_{N_{p}}\right)$ and $p_{x,N_{p}+i}$. Due to the proportionality between $f\left(z_{N_{p}+i}|\hat{z}_{N_{p}+i},\sigma_{N_{p}+i}\right)$ and $f\left(x_{N_{p}+i}|x_{1},x_{2},\cdots,x_{N_{p}}\right)$ (see section 3.3), we can therefore conclude that the relationship between $p_{z,N_{p}+i}$ and $p_{x,N_{p}+i}$ is also strictly monotonically increasing. This indicates that the ordering of values of $p_{x,N_{p}+i}$ is preserved in its proxy $p_{z,N_{p}+i}$, making this estimated probability of conformance in the latent space a useful metric for anomaly detection.
@@ 4.1 Proportionality test
The pipeline consists of an encoder with a single fully connected layer to a latent space encoding with a size of $N_{e}=4$ and a decoder with a fully connected layer back to a single value. A gated recurrent unit (GRU) model with $N_{g}=8$ units is used for predictions together with a forecasting model consisting of six fully connected layers with $[16,16,32,32,64,64]$ units, followed by two separate fully connected layers for the estimation of $\hat{z}_{2}$ (linear activation) and $\sigma_{2}$ (exponential activation). Training is performed using the RMSprop optimizer ($lr=1\cdot 10^{-4}$, $\rho=0.9$) with a batch size of 64. The models are optimized for 1000 warm-up iterations based on the loss in Equation 4, after which training with the normal loss function (Equation 3) continues until convergence. The loss weight for the reconstruction loss is set to $\lambda=100$. Training is repeated 100 times to determine sensitivity to initial conditions, resulting in 100 models.
@@ 4.2 Sine wave frequency deviation
A signal containing the instantaneous frequencies at every time step $f(t)$ is generated using a constrained random walk, such that $f_{c}(t)-\frac{f_{b}}{2}\leq f(t)\leq f_{c}(t)+\frac{f_{b}}{2}$, where $f_{b}=0.25$ Hz is the frequency band. Furthermore, bounded random walks were used to create a baseline wander $1\leq b(t)\leq 1$ and wave amplitude $0.5\leq a(t)\leq 2$. Lastly, additive noise was $n(t)\sim\mathcal{N}\left(0,\sigma_{n}^{2}\right)$, where $\sigma_{n}\sim\mathcal{U}(0,0.2)$. The corresponding sine wave signal was then determined by
$s(t)=a(t)\cdot\sin\left(2\pi\int_{0}^{t}f(t)dt\right)+b(t)+n(t).$ (11)
@@
Using this data model, we synthesized a training and validation dataset of 1,000,000 and 100,000 signals respectively at a sampling frequency of $f_{s}=128$ Hz. The signals consist of 8 segments of 256 samples each, totaling 2048 samples (16 seconds). The first $5$ segments $x_{1}$ to $x_{5}$ are used to predict the latent space representations of the last $3$ segments $x_{6}$ to $x_{8}$. The center frequencies were set such that $f_{c,before}=f_{c,after}$, implying the absence of change points. In Figure 3, examples of normal and abnormal signals are shown.
@@
Figure 3: Examples of generated signals. The top signal has no frequency deviation. The center signal has a frequency change of 11 Hz at the change point around $t=11$ s. The bottom signal has a frequency change of 22 Hz at the change point around $t=11$ s.
We used an encoder model $E$ with two blocks (convolutional, ReLU, BatchNorm, Maxpooling) with an increasing amount of filters (32 and 64), followed by a fully connected layer to a latent space with size $N_{e}=16$. The accompanying decoder $D$ follows a similar structure with a fully connected layer, followed by two blocks (convolutional, ReLU, upsampling) and one final convolutional layer. As the recurrent neural network $G$, we used a GRU with 32 units. The three forecasting models $F_{1}$, $F_{2}$, and $F_{3}$ consist of three fully connected layers (64, 128, and 256 units), followed by two separate fully connected layers for the estimation of $\hat{z}_{6}$ to $\hat{z}_{8}$ (linear activation) and $\sigma_{6}$ to $\sigma_{8}$ (exponential activation).
All models were simultaneously optimized on an Intel Core i9-10980XE CPU and NVIDIA RTX A6000 GPU hardware set-up with Windows Server 2022, Python 3.9.15 and TensorFlow 2.10. An RMSprop optimizer with a learning rate of $1\cdot 10^{-4}$ and a batch size of $32$ was used. Training consisted of 1000 warm-up iterations ($0.89$ seconds/iteration), followed by normal training until convergence on the validation dataset ($\sim 365,000$ iterations, $0.58$ seconds/iteration). The reconstruction loss weight was set to $\lambda=10^{4}$. The model optimization required $\sim 1800$ MB of VRAM.
After training, classification performance was tested on two test datasets. Both were generated in a similar manner as the training and validation set, except that change points could be introduced in segment $x_{6}$. Both datasets consisted of $100,000$ signals with anomalous change points and $100,000$ without anomalous change points. The first test dataset was used to determine an optimal classification threshold that maximizes the F1 score. The second test dataset, together with the determined classification threshold, was used to determine several classification performance metrics, which are provided in Table 2. Receiver Operating Characteristic (ROC) and Precision-Recall (PR) curves are shown in Figure 4. The classification performance metrics, as well as the ROC/PR curves, suggest that the PPC pipeline has a strong discriminative ability. Furthermore, inference time and memory were measured on the same machine, resulting in $0.098$ seconds/sample and $\sim 1200$ MB of VRAM.
Table 2: Anomalous change point classification performance metrics for the Sine wave frequency deviation experiment (section 4.2) and Counting with MNIST digits experiment (section 4.3).
Metric
PPC-SINE
(Section 4.2)
PPC-MNIST
(Section 4.3)
ROC AUC $0.989$ $0.946$
PR AUC $0.992$ $0.993$
TP (True Positives) $95910$ $88347875$
FP (False Positives) $685$ $3925349$
TN (True Negatives) $99315$ $6080573$
FN (False Negatives) $4090$ $1646203$
Recall $95.9\%$ $98.2\%$
Precision $99.3\%$ $95.7\%$
Specificity $99.3\%$ $60.8\%$
Balanced accuracy $97.6\%$ $79.5\%$
Matthews Correlation Coefficient $0.953$ $0.662$
F1 Score $0.976$ $0.969$
[…]
@@
Figure 4: Receiver Operating Characteristic and Precision-Recall curves for the application of the PPC pipeline on the sine wave frequency deviation example.
Furthermore, to visualize the response to different pairs of $f_{c,before}$ and $f_{c,after}$, we synthesized 10 signals for every combination of these center frequencies, with a resolution of $0.05$ Hz. This results in $\left(\frac{10-0.5}{0.05}\right)^{2}\cdot 10=361,000$ signals. We applied the model to these signals and calculated the log-likelihood and probability of conformance for every pair of center frequencies $\left(f_{c,before},f_{c,after}\right)$. These metrics were averaged and are shown in Figure 5, together with the frequency band with which the training and validation signals were generated. As expected, the probability of conformance is high if $f_{c,before}\approx f_{c,after}$, and low for other combinations of center frequencies. This response suggests that the PPC pipeline is able to detect anomalous change points over a wide range of center frequencies.
Figure 5: Average log-likelihood (left) and probability of conformance (right) per combination of center frequencies. The frequency band margins indicate how much frequencies may fluctuate as part of the generated instantaneous frequencies $f(t)$.
@@ 4.3 Counting with MNIST digits
In this experiment, we used the Modified National Institute of Standards and Technology (MNIST) dataset [44], consisting of 60,000 training images and 10,000 test images of handwritten digits. The objective within this experiment for the pipeline is to determine whether the corresponding digits of a set of two MNIST images are sequential, e.g., if $x_{1}$ formed an image of the digit 3, $x_{2}$ should represent the digit 4. In the special case of digit 9, we considered 0 to be its successor. Examples of sequences that (do not) follow this pattern are provided in Figure 6. The experiment was designed to determine if the pipeline is able to capture (deviations from) patterns in relatively complex sequential image data.
Figure 6: Examples of correct (green) and incorrect (red) sequences of MNIST digits. During training, only correct sequences of digits are used.
@@
First, 10,000 validation images were selected from the original set of training images, reflecting the class composition of the training set. All image pixels were normalized to the range $[0,1]$. Data augmentation is applied online, consisting of random rotation (between $18^{\circ}$ and $18^{\circ}$), random rectangular zooming (between $5\%$ and $5\%$), and random translation (between $5\%$ and $5\%$).
@@
We used an encoder model $E$ with two blocks (convolutional, ReLU, BatchNorm, Maxpooling) with an increasing amount of filters (32 and 64), followed by a fully connected layer to a latent space with size $N_{e}=16$. The accompanying decoder $D$ follows a similar structure with a fully connected layer, followed by two blocks (convolutional, ReLU, upsampling) and one final convolutional layer. As the recurrent neural network $G$, we used a GRU with 32 units. The single forecasting model $F_{1}$ consists of three fully connected layers (64, 128, and 256 units), followed by two separate fully connected layers for the estimation of $\hat{z}_{2}$ (linear activation) and $\sigma_{2}$ (exponential activation).
All models were simultaneously optimized on an Intel Core i9-10980XE CPU and NVIDIA RTX A6000 GPU hardware set-up with Windows Server 2022, Python 3.9.15 and TensorFlow 2.10. An RMSprop optimizer with a learning rate of $1\cdot 10^{-4}$ and a batch size of $32$ was used. Training consisted of 1000 warm-up iterations ($0.57$ seconds/iteration), followed by normal training until convergence ($\sim 132,000$ iterations, $0.57$ seconds/iteration. The reconstruction loss weight was set to $\lambda=10^{4}$. The model optimization required $\sim 1300$ MB of VRAM.
After training, the pipeline was used to compare every pair of two images in the test set, yielding a log-likelihood and probability of conformance for every pair. Similarly to the Sine wave frequency deviation experiment (Section 4.2), a classification threshold was determined on the validation set, and classification performance metrics were calculated based on the test set. These metrics are reported in Table 2 and suggest strong discriminative ability of the PPC pipeline. ROC and PR curves are shown in Figure 8. Inference took $0.026$ seconds/sample and $\sim 1100$ MB of VRAM.
@@
To see the pipeline’s response to different combinations of digits, log-likelihood and probability of conformance were averaged over every pair of digits and are shown in Figure 7. As expected, the log-likelihoods and probabilities of conformance on the subdiagonal and top right are significantly higher than for other combinations of digits. This indicates that the model indeed is able to differentiate between successive digits and non-successive digits. Some combinations, like the combination for $(3,9)$, show higher values due to the similarity between digits, e.g., a handwritten 4 is similar in shape to the digit 9.
@@
Figure 7: Average log-likelihood (left) and probability of conformance (right) per combination of digits.
@@
Figure 8: Receiver Operating Characteristic and Precision-Recall curves for the application of the PPC pipeline on the successive MNIST digits example.
@@ 4.4 Artifact detection for MRSI
After training, the pipeline was used to predict all segments of every spectrum of an example in-vivo 2D MRSI slice. We further calculate the likelihoods and probabilities of conformance for every segment. Figure 10 depicts the probabilities of conformance for six example spectra located at the indicated locations as well as a probability map with pixels representing averaged spectra. Moving from the top left to top right spectrum: spectra within the homogeneous regions of the brain appear normal and show the highest probabilities of conformance, measurements near the ventricles are distorted completely, and susceptibility artifacts coming from the nasal cavity are present and detected by the method. From bottom left to right: non-anomalous spectra are consistently labeled with high probabilities, spectra close to the skull show residual lipid signal artifacts and lower probabilities of conformance, and the spectra outside the brain mask are nulled leading to no anomalies.
Figure 10: In-vivo 2D MRSI slice with indicated locations of selected segmented example spectra. The probability of conformance is shown for each segment of the spectra as well as averaged for each pixel (representing a single voxel).
@@
Figure 11: 2D brain maps for the estimated likelihoods, the metabolite concentration map for total N-acetylaspartate (NAA + NAAG) estimated by LCModel [70], as well as the corresponding CRLB% (left to right).
@@ 5 Discussion
In contrast to some other methods mentioned in section 2, anomaly detection with the proposed PPC method has a linear time complexity ($\mathcal{O}(n)$) with respect to the number of samples $n$, as individual samples are processed independently. This allows for scalability to large datasets. While training the PPC pipeline can be computationally expensive compared to non-neural network approaches, the models used in our experiments are remarkably lightweight compared to other deep neural network architectures. The models used in sections 4.2, 4.3, and 4.4 have $5.8\cdot 10^{5}$, $2.3\cdot 10^{5}$ and $9.5\cdot 10^{5}$ trainable parameters, respectively, while modern deep neural networks have significantly more parameters. For example, Inceptionv3 [77], ResNet-50 [30], and VGG-16 [73] have about $2.4\cdot 10^{7}$, $2.5\cdot 10^{7}$, and $1.4\cdot 10^{8}$ trainable parameters, respectively. Furthermore, where other methods may fail to provide meaningful or intuitive anomaly scores, PPC is able to provide an anomaly score that is not only normalized to a closed unit interval $[0,1]$, but also indicates a probability of data conforming to what was expected.
@@
The artifact detection for MRSI experiment (section 4.4) shows the real-world application of the method to MRSI data. The probability of conformance delivers a valuable measure for the quality of MRSI spectra and enables the filtering of artifact-ridden spectra by setting a specific threshold. This is particularly useful with high‐resolution whole‐brain acquisition techniques where hundreds of thousands of spectra are acquired while only good-quality spectra can be quantified reliably [54, 56]. Note the CRLB% is only obtained after computationally intensive least-squares fitting of the spectra. In addition, having a segment-wise quality measure for each spectrum can allow metabolite-targeted filtering to improve studies of specific brain pathologies [9, 26, 68].
@@
Lastly, we recognize the lack of benchmark evaluations in this work. The lack of available real-world datasets forms an obstacle to the development and testing of CPD and AD methods [64]. This is even more problematic for ACPD methods such as PPC. All works mentioned in section 2.4 use specifically-designed synthetic datasets, private datasets, datasets containing only normal data and anomalies (for testing AD methods), or datasets containing only normal data and abnormal change points (for CPD testing methods). Furthermore, these datasets are often low-complexity, univariate time series data that are outdated and/or unannotated. Therefore, we strongly recommend that future research on ACPD methods includes the development of publicly available benchmark datasets that form the standard for the evaluation of ACPD methods.